[Brahms-l] Shift report 20050424 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Apr 24 2005 - 01:15:33 EDT
Shifter leader was kh
Shifters were 


Had beam during most of the shift.
Runs achieved
Run   tr2   tr3
14226 ZDC Min bias run (637119 tr4)
14228  6726  132262
14229  5446  111136
14233  5240   50794
14234  ZDC min bias run (140420 tr4)

Problems:
Mainframe BBR1 which serves CCR1-5, MRST0 and BBR1-15 started to trip.  This may have been going on since last night.  Problem was discovered about 16:30.  We got temporary relief by turning of BBR1-15; that apparently enabled the CC and MRST0 counters to operate.  By 23:00, even that could not be sustained.  B. Scheetz is in and we have a 15 min access to try to install a spare LR4032a

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Apr 24 01:15:48 2005

This archive was generated by hypermail 2.1.8 : Sun Apr 24 2005 - 01:16:07 EDT